Your company just paid a dividend of $3. Management of your company wants to expand with a new product line. This would cause your company’s constant growth rate in earnings and dividends to rise from 5% to 9%. Currently, the require rate of return for your company is 10%. This expansion would cause the required rate of return to increase to 13%, due to additional risk. Would shareholders want management to make this change? (Hint, P0). Show calculations to support your YES or NO answer for full credit.
Answer #1

Using DDM Model,

Stock Price without new Product line = 3(1.05)/(0.10 - 0.05) = $63

Stock Price with new Product line = 3(1.09)/(0.13 - 0.09) = $81.75

As stock Price is increasing after new product line, so shareholders will want this new investment.